OpenStudy (anonymous):

Water coming out from a fountain is modeled by the function f(x)=-x^2+8x+2 where f(x) represents the height, in feet, of the water from the fountain at different times x, in seconds. What does the average rate of change f(x) from x=1 to x=4 represent? The water travels an average distance of 3 feet from 1 second to 4 seconds. The water travels an average distance of 2 feet from 1 second to 4 seconds. The water rises up with an average speed of 3 feet per second from 1 second to 4 seconds. The water rises up with an average speed of 2 feet per second from 1 second to 4 seconds.

OpenStudy (anonymous):

Please explain how instead of just saying the answer, thanks!

OpenStudy (anonymous):

@SamsungFanBoy

OpenStudy (anonymous):

@Dolkun

OpenStudy (anonymous):

@MikeyMaximum

OpenStudy (anonymous):

@mrbaconqwerty

OpenStudy (mrbaconqwerty):

Dont know. Sorry :(

OpenStudy (anonymous):

That's fine, thank you for at least checking it out (:

OpenStudy (anonymous):

@Chris01 ?

OpenStudy (anonymous):

@Tuta

OpenStudy (anonymous):

I feel like it's either A or C but I'm not sure how to differentiate (sorry if I spelt that wrong) the two

OpenStudy (anonymous):

don't really know how to explain it. but im pretty sure its B

OpenStudy (anonymous):

Do you know how you got b? :/

OpenStudy (anonymous):

find the values of f at x = 1 and x = 4. Then you'll subtract these values. This will give you the absoulte change. Then divide the absolute change over the interval by the length of the interval, to find the average amount of change in f(x) per one-unit change in x

OpenStudy (anonymous):

I did that, and I got (1,9) and (4,18) I subtracted 18 by 9 and got 9, and 4-1=3. so 9/3=3 and that's why I thought it was A or C

OpenStudy (anonymous):

can you help figure this out @Ashleyisakitty

OpenStudy (anonymous):

cause i'm still thinking its B

OpenStudy (anonymous):

Can you maybe input the numbers and everything so I know how you got B?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

I'm leaning toward C and D now /:

OpenStudy (anonymous):

i'm lagging so bad sorry if i'm taking so long

OpenStudy (anonymous):

It's totally fine, since I do online, I could turn this in whenever. Haha.

OpenStudy (anonymous):

I just don't want to be wrong because this is on one of my final exams and yeah :|

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

@corruptedalice96 The first two choices said "the average distance travelled" which implies the curved distance travelled. The question asks for the change in f(x), which relates to the height. This information together with Chris01's will enable you to get the right answer.

OpenStudy (anonymous):

I'm still confused, do you think you could explain, but with different numbers?

OpenStudy (anonymous):

I'm pretty confident it's C but I don't know if that's right

OpenStudy (anonymous):

@mathmate ?

OpenStudy (mathmate):

f(x)=-x^2+8x+2 differentiate f(x) with respect to x: f'(x)=-2x+8 where f'(x) represents the vertical speed (rate of change) as a function of time. f'(1)=+6 f'(4)=+0 Average speed (rate of change) = (f'(1)+f'(4))/2 = (6+0)/2 = 3 fps To find the derivative (i.e. result of differentiation) see http://www.wikihow.com/Differentiate-Polynomials

OpenStudy (anonymous):

I'm so lost..

OpenStudy (mathmate):

 the drawings should help. If you would review this term's work on derivatives and functions, it would be much easier to follow.

OpenStudy (anonymous):

@mathmate I'm going with C then..

OpenStudy (anonymous):

@♪Chibiterasu do you think you could check my answer?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

ITS THE ANSWER C OR D? @CORRUPTEDALICE96